# Protrusion Decompositions Revisited: Uniform Lossy Kernels for Reducing Treewidth and Linear Kernels for Hitting Disconnected Minors

### v1

@@ Abstract
In this work, we show that this non-uniformity can be avoided at the expense of a small loss. First, we present a simple 2-approximate kernelization algorithm for Treewidth-$\eta$-Deletion with kernel size $g(\eta)\cdot k^{5}$. Next, we show that the approximation factor can be made arbitrarily close to $1$, if we settle for a kernelization protocol with $\mathcal{O}(1)$ calls to an oracle that solves instances of size bounded by a uniform polynomial in $k$. We extend the above results to general $\mathcal{F}$-Deletion, whenever $\mathcal{F}$ contains a planar graph, as long as an oracle for Treewidth-$\eta$-Deletion is available for small instances. Notably, all our constants are computable functions of $\mathcal{F}$ and our techniques work also when some graphs in $\mathcal{F}$ may be disconnected.
@@ 1 Introduction
What may seem like a downside of the kernelization algorithm by Fomin et al. [DBLP:conf/focs/FominLMS12] is its non-uniformity: the size of the kernel is of the form $\mathcal{O}_{\mathcal{F}}(k^{f(\mathcal{F})})$ for some function $f$. This turns out to be inevitable: Treewidth-$\eta$-Deletion does not admit a kernel of size $\mathcal{O}_{\eta}(k^{o(\eta)})$ unless NP $\subseteq$ coNP/poly [DBLP:journals/talg/GiannopoulouJLS17]. On the positive side, uniform kernels are known for Treedepth-$\eta$-Deletion with size $\mathcal{O}_{\eta}(k^{6})$ [DBLP:journals/talg/GiannopoulouJLS17] and for families $\mathcal{F}$ of connected graphs that include $K_{2,p}$ (for any $p\in\mathbb{N}$) with size $\mathcal{O}_{\mathcal{F}}(k^{10})$ [DBLP:conf/isaac/LochetS24] (cf. [DBLP:journals/siamdm/FominLMPS16]).
@@
Let $\mathcal{F}$ be a finite family of graphs containing at least one planar graph. Then $\mathcal{F}$-Deletion admits a $2$-lossy compression of size $\mathcal{O}_{\mathcal{F}}(k^{5})$. Moreover, Treewidth-$\eta$-Deletion admits a $2$-lossy kernel of size $\mathcal{O}_{\eta}(k^{5})$.
@@
Let $\mathcal{F}$ be a finite family of graphs containing at least one planar graph. For any $\epsilon>0$, $\mathcal{F}$-Deletion admits a $(1+\epsilon)$-lossy compression protocol of call size $\mathcal{O}_{\mathcal{F}}(k^{5}+k^{3r})$, and at most $1+r$ rounds, where $r=1+\left\lceil\log_{1+\epsilon}(\frac{1}{\epsilon})\right\rceil$.
@@ 2 Preliminaries
For any object $F$, the notation $\mathcal{O}_{F}(\cdot)$ hides factors depending on $F$. For $p\in\mathbb{N}$ we denote $p]=[1,p]=\{1,\ldots,p\}$. By the size of a graph, we mean the number of vertices in it.
@@
A $t$-boundaried graph is a triple ${\bf H}=(H,B,\lambda)$ where $H$ is a graph, $B\subseteq V(H)$ is of size $t$ and $\lambda\colon[t]\to B$ is a bijection representing a labeling function. We refer to the set $B$ as $\partial{\bf H}$ and call it the boundary of $\bf H$}. By $V({\bf H})$ we refer to $V(H)$. Given two $t$-boundaried graphs ${\bf H_{1}}=(H_{1},B_{1},\lambda_{1})$ and ${\bf H_{2}}=(H_{2},B_{2},\lambda_{2})$, ${\bf H_{1}}\oplus{\bf H_{2}}$ denotes the graph $H$ obtained by first taking the disjoint union of $H_{1}$ and $H_{2}$ and then identifying each vertex $u_{1}$ in $B_{1}$ with $u_{2}\in B_{2}$ such that $\lambda_{1}^{-1}(u_{1})=\lambda_{2}^{-1}(u_{2})$. For any $h\in\mathbb{N}$ and a graph $G$, $h$-folio(G)(G) is the set of all graphs on at most $h$ vertices which appear in $G$ as a minor.
@@ 3 Technical overview
We would like to transform the current decomposition into a $(\mathcal{O}_{\eta}(k^{c}),\mathcal{O}_{\eta}(1))$-protrusion decomposition, where $c$ does not depend on $\eta$. We call a component $C$ of $G-(X\cup Z)$ simplicial if $N(C)$ forms a clique. Because we can insert edges between highly connected vertices in $X\cup Z$, each non-edge may appear only in a bounded number of sets $N(C)$. We use this argument to bound the number of non-simplicial components by $\mathcal{O}_{\eta}(k^{7})$ (this bound can further improved by making a finer analysis which we postpone to the place where the formal proof appears). Let us neglect the simplicial components for a moment and let $G^{\prime}$ denote the graph without them. Then $G^{\prime}$ admits a protrusion decomposition of desired form and we can compress $G^{\prime}$ to $G^{\prime\prime}$ of size $\mathcal{O}_{\eta}(k^{7})$. This constitutes the output of the kernelization algorithm.
@@
Suppose first that $|Y_{2}|>(1-\varepsilon)\cdot|Y_{1}|$. Observe that any optimal solution $S$ in $G$ satisfies $|S\cap Y_{1}|\geq\mathsf{opt}(G[Y_{1}])$ so $\mathsf{opt}(G-Y_{1})\leq\mathsf{opt}(G)-(1-\varepsilon)\cdot|Y_{1}|$. A crucial observation is that removing $Y_{1}$ from $G$ allows us to construct an $(\mathcal{O}_{\eta}(k^{7}),\mathcal{O}_{\eta}(1))$-protrusion decomposition. For each simplicial component $C$, the set $N(C)\setminus Y_{1}$ forms a clique in the graph $G[Y_{0}\setminus Y_{1}]$ of treewidth $\leq\eta$. But the number of cliques in a bounded-treewidth graph is linear so we can group the simplicial components into $\mathcal{O}_{\eta}(|Y_{0}\setminus Y_{1}|)$ protrusions according to their neighborhoods in $Y_{0}\setminus Y_{1}$. Together with $\mathcal{O}_{\eta}(k^{7})$ non-simplicial components, this allows us to compress the graph $G-Y_{1}$ into size $\mathcal{O}_{\eta}(k^{7})$. Let $S_{1}$ be the solution in $G-Y_{1}$ computed with the help of the oracle. Assuming $|S_{1}|=\mathsf{opt}(G-Y_{1})$ we get $|S_{1}|+|Y_{1}|\leq\mathsf{opt}(G)-(1-\varepsilon)\cdot|Y_{1}|+|Y_{1}|=\mathsf{opt}(G)+\epsilon\cdot|Y_{1}|\leq(1+\epsilon)\cdot\mathsf{opt}(G)$.
Suppose now that $|Y_{2}|\leq(1-\varepsilon)\cdot|Y_{1}|$. We continue asking the oracle for solution $Y_{i+1}$ in $G[Y_{i}]$. Let $j$ be the first iteration for which $|Y_{j+1}|>(1-\varepsilon)\cdot|Y_{j}|$. We repeat the same trick but this time the set $Y_{0}\setminus Y_{j}$ is partitioned into $j$ subsets $(Y_{0}\setminus Y_{1}),(Y_{1}\setminus Y_{2}),\dots,(Y_{j-1}\setminus Y_{j})$, each inducing a graph of treewidth $\leq\eta$. The number of cliques in $G[Y_{0}\setminus Y_{j}]$ can be bounded by $f(\eta,j)\cdot|Y_{0}|^{j}$, allowing us to construct an $(\mathcal{O}_{\eta,j}(k^{\mathcal{O}(j)+7}),\mathcal{O}_{\eta,j}(1))$-protrusion decomposition.
@@ 5 Uniform $2$-lossy polynomial compression algorithm
Let $\mathcal{P}_{\textnormal{{non-simp}}}$ be the set of non-simplicial parts of $(P_{1},\ldots,P_{\ell})$. From Lemma 5.5, $|\mathcal{P}_{\textnormal{{non-simp}}}|=\mathcal{O}_{\eta}(k^{7})$. Run the reduction algorithm of ˜2 on the instance $({G^{1}}_{\textnormal{{flow}}}[V(G^{1})\setminus\bigcup_{P_{i}\in\mathcal{P}_{\textnormal{{simp}}}}P_{i}],k_{1})$ of $\mathcal{F}$-Deletion. It returns an instance $(G^{2},k_{1})$ of Treewidth-$\eta$-Deletion such that $G^{2}$ has an $(\mathcal{O}_{\eta}(k^{7}),\mathcal{O}(\eta))$-protrusion decomposition. Let $(G^{3},k_{1})$ be the instance returned by the algorithm of Lemma 4.6 on input $(G^{2},k_{1})$. Observe that $k_{1}\leq k$. From Lemma 4.6 the size of $G^{3}$ is bounded by $\mathcal{O}_{\mathcal{F}}(k^{7})$.
@@ 6 Uniform $(1+\epsilon)$-lossy polynomial compression protocol
Fix any $\epsilon>0$. We describe a polynomial time $(1+\epsilon)$-lossy compression protocol of capacity $\mathcal{O}_{\eta}(k^{7})+\mathcal{O}_{\eta}(k^{3r})$ and $1+r$ rounds. For any $\beta\geq 1$ it has access to $\beta$-kernelization oracles for $\mathcal{F}$-Deletion and Treewidth-$\eta$-Deletion of capacities $\mathcal{O}_{\eta}(k^{7})+\mathcal{O}_{\eta}(k^{3r})$.
@@
We obtain that $G^{2}$ has a $(\delta,\gamma)$-protrusion decomposition, where $\delta=\mathcal{O}_{\eta}(k^{7})+\mathcal{O}_{\eta}(k^{3r})$ and $\gamma=\mathcal{O}(\eta/\epsilon)$. Let $(G^{3},k_{3})$ be the instance obtained from Lemma 4.6 on input $(G^{2},k_{2})$. Then the size of $G^{3}$ is $\mathcal{O}_{\mathcal{F}}(\delta)$. Finally run the $\beta$-kernelization algorithm for the problem $\mathcal{F}$-Deletion on the instance $(G^{3},k_{3})$. Let $S_{3}$ be a $\beta$-approximate solution for the problem $\mathcal{F}$-Deletion on the instance $(G^{3},k_{3})$.

### v2

@@ Abstract
In this work, we show that this non-uniformity can be avoided at the expense of a small loss. First, we present a simple 2-approximate kernelization algorithm for Treewidth-$\eta$-Deletion with a kernel of size $g(\eta)\cdot k^{6}$. Next, we show that the approximation factor can be made arbitrarily close to $1$, if we settle for a kernelization protocol with $\mathcal{O}(1)$ calls to an oracle that solves instances of size bounded by a uniform polynomial in $k$. We extend the above results to general $\mathcal{F}$-Deletion, whenever $\mathcal{F}$ contains a planar graph, as long as an oracle for Treewidth-$\eta$-Deletion is available for small instances. Notably, all our constants are computable functions of $\mathcal{F}$ and our techniques work also when some graphs in $\mathcal{F}$ may be disconnected.
@@ 1 Introduction
What may seem like a downside of the kernelization algorithm by Fomin et al. [DBLP:conf/focs/FominLMS12] is its non-uniformity: the size of the kernel is of the form $\mathcal{O}_{\mathcal{F}}(k^{f(\mathcal{F})})$ for some function $f$. This turns out to be inevitable: Treewidth-$\eta$-Deletion does not admit a kernel of size $\mathcal{O}_{\eta}(k^{o(\eta)})$ unless NP $\subseteq$ coNP/poly [DBLP:journals/talg/GiannopoulouJLS17]. On the positive side, uniform kernels are known for Treedepth-$\eta$-Deletion with $\mathcal{O}_{\eta}(k^{6})$ vertices [DBLP:journals/talg/GiannopoulouJLS17] and for families $\mathcal{F}$ of connected graphs that include $K_{2,p}$ (for any $p\in\mathbb{N}$) with $\mathcal{O}_{\mathcal{F}}(k^{10})$ vertices [DBLP:conf/isaac/LochetS24] (cf. [DBLP:journals/siamdm/FominLMPS16]).
@@
Let $\mathcal{F}$ be a finite family of graphs containing at least one planar graph. Then $\mathcal{F}$-Deletion admits a $2$-lossy compression with $\mathcal{O}_{\mathcal{F}}(k^{5})$ vertices and of size $\mathcal{O}_{\mathcal{F}}(k^{6})$. Moreover, Treewidth-$\eta$-Deletion admits a $2$-lossy kernel with $\mathcal{O}_{\mathcal{F}}(k^{5})$ vertices and of size $\mathcal{O}_{\mathcal{F}}(k^{6})$.
@@
Let $\mathcal{F}$ be a finite family of graphs containing at least one planar graph. For any $\epsilon>0$, $\mathcal{F}$-Deletion admits a $(1+\epsilon)$-lossy compression protocol of call size $\mathcal{O}_{\mathcal{F}}(k^{6}+k^{3r+1})$, and at most $1+r$ rounds, where $r=1+\left\lceil\log_{1+\epsilon}(\frac{1}{\epsilon})\right\rceil$.
@@ 2 Preliminaries
For any object $F$, the notation $\mathcal{O}_{F}(\cdot)$ hides factors depending on $F$. For $p\in\mathbb{N}$ we denote $p]=[1,p]=\{1,\ldots,p\}$.
@@
A $t$-boundaried graph is a triple ${\bf H}=(H,B,\lambda)$ where $H$ is a graph, $B\subseteq V(H)$ is of size $t$ and $\lambda\colon[t]\to B$ is a bijection representing a labeling function. We refer to the set $B$ as $\partial{\bf H}$ and call it the boundary of $\bf H$}. By $V({\bf H})$ we refer to $V(H)$. Given two $t$-boundaried graphs ${\bf H_{1}}=(H_{1},B_{1},\lambda_{1})$ and ${\bf H_{2}}=(H_{2},B_{2},\lambda_{2})$, ${\bf H_{1}}\oplus{\bf H_{2}}$ denotes the graph $H$ obtained by first taking the disjoint union of $H_{1}$ and $H_{2}$ and then identifying each vertex $u_{1}$ in $B_{1}$ with $u_{2}\in B_{2}$ such that $\lambda_{1}^{-1}(u_{1})=\lambda_{2}^{-1}(u_{2})$.
For any $h\in\mathbb{N}$ and a graph $G$, $h$-folio(G)(G) is the set of all graphs on at most $h$ vertices which appear in $G$ as a minor.
@@ 3 Technical overview
We would like to transform the current decomposition into a $(\mathcal{O}_{\eta}(k^{c}),\mathcal{O}_{\eta}(1))$-protrusion decomposition, where $c$ does not depend on $\eta$. We call a component $C$ of $G-(X\cup Z)$ simplicial if $N(C)$ forms a clique. Because we can insert edges between highly connected vertices in $X\cup Z$, each non-edge may appear only in a bounded number of sets $N(C)$. We use this argument to bound the number of non-simplicial components by $\mathcal{O}_{\eta}(k^{5})$. Let us neglect the simplicial components for a moment and let $G^{\prime}$ denote the graph without them. Then $G^{\prime}$ admits a protrusion decomposition of desired form and we can compress $G^{\prime}$ to $G^{\prime\prime}$ on $\mathcal{O}_{\eta}(k^{5})$ vertices. This constitutes the output of the kernelization algorithm.
@@
Suppose first that $|Y_{2}|>(1-\varepsilon)\cdot|Y_{1}|$. Observe that any optimal solution $S$ in $G$ satisfies $|S\cap Y_{1}|\geq\mathsf{opt}(G[Y_{1}])$ so $\mathsf{opt}(G-Y_{1})\leq\mathsf{opt}(G)-(1-\varepsilon)\cdot|Y_{1}|$. A crucial observation is that removing $Y_{1}$ from $G$ allows us to construct an $(\mathcal{O}_{\eta}(k^{7}),\mathcal{O}_{\eta}(1))$-protrusion decomposition. For each simplicial component $C$, the set $N(C)\setminus Y_{1}$ forms a clique in the graph $G[Y_{0}\setminus Y_{1}]$ of treewidth $\leq\eta$. But the number of cliques in a bounded-treewidth graph is linear so we can group the simplicial components into $\mathcal{O}_{\eta}(|Y_{0}\setminus Y_{1}|)$ protrusions according to their neighborhoods in $Y_{0}\setminus Y_{1}$. Together with $\mathcal{O}_{\eta}(k^{5})$ non-simplicial components, this allows us to compress the graph $G-Y_{1}$ into size $\mathcal{O}_{\eta}(k^{5})$. Let $S_{1}$ be the solution in $G-Y_{1}$ computed with the help of the oracle. Assuming $|S_{1}|=\mathsf{opt}(G-Y_{1})$ we get $|S_{1}|+|Y_{1}|\leq\mathsf{opt}(G)-(1-\varepsilon)\cdot|Y_{1}|+|Y_{1}|=\mathsf{opt}(G)+\epsilon\cdot|Y_{1}|\leq(1+\epsilon)\cdot\mathsf{opt}(G)$.
Suppose now that $|Y_{2}|\leq(1-\varepsilon)\cdot|Y_{1}|$. We continue asking the oracle for solution $Y_{i+1}$ in $G[Y_{i}]$. Let $j$ be the first iteration for which $|Y_{j+1}|>(1-\varepsilon)\cdot|Y_{j}|$. We repeat the same trick but this time the set $Y_{0}\setminus Y_{j}$ is partitioned into $j$ subsets $(Y_{0}\setminus Y_{1}),(Y_{1}\setminus Y_{2}),\dots,(Y_{j-1}\setminus Y_{j})$, each inducing a graph of treewidth $\leq\eta$. The number of cliques in $G[Y_{0}\setminus Y_{j}]$ can be bounded by $f(\eta,j)\cdot|Y_{0}|^{j}$, allowing us to construct an $(\mathcal{O}_{\eta,j}(k^{\mathcal{O}(j)+5}),\mathcal{O}_{\eta,j}(1))$-protrusion decomposition.
@@ 5 Uniform $2$-lossy polynomial compression algorithm
Let $\mathcal{P}_{\textnormal{{non-simp}}}$ be the set of non-simplicial parts of $(P_{1},\ldots,P_{\ell})$. From Lemma 5.5, $|\mathcal{P}_{\textnormal{{non-simp}}}|=\mathcal{O}_{\eta}(k^{5})$. Run the reduction algorithm of ˜2 on the instance $({G^{1}}_{\textnormal{{flow}}}[V(G^{1})\setminus\bigcup_{P_{i}\in\mathcal{P}_{\textnormal{{simp}}}}P_{i}],k_{1})$ of $\mathcal{F}$-Deletion. It returns an instance $(G^{2},k_{1})$ of Treewidth-$\eta$-Deletion such that $G^{2}$ has an $(\mathcal{O}_{\eta}(k^{5}),\mathcal{O}(\eta))$-protrusion decomposition. Let $(G^{3},k_{1})$ be the instance returned by the algorithm of Lemma 4.6 on input $(G^{2},k_{1})$. Observe that $k_{1}\leq k$. From Lemma 4.6 the number of vertices of $G^{3}$ is bounded by $\mathcal{O}_{\mathcal{F}}(k^{5})$. Observe that if $G^{3}$ indeed has an at most $k$-sized vertex set whose deletion results in a graph of treewidth at most $\eta$, then treewidth of $G^{3}$ has to be at most $k+\eta$. Thus, the number of edges of $G^{3}$, and hence the size of $G^{3}$, is at most $(k+\eta)$ times the number of vertices of $G^{3}$. Thus, the size of $G^{3}$ is $\mathcal{O}_{\mathcal{F}}(k^{6})$, otherwise $G^{3}$ has no $k$-sized solution.
@@ 6 Uniform $(1+\epsilon)$-lossy polynomial compression protocol
Fix any $\epsilon>0$. We describe a polynomial time $(1+\epsilon)$-lossy compression protocol of capacity $\mathcal{O}_{\eta}(k^{7})+\mathcal{O}_{\eta}(k^{3r})$ and $1+r$ rounds. For any $\beta\geq 1$ it has access to $\beta$-kernelization oracles for $\mathcal{F}$-Deletion and Treewidth-$\eta$-Deletion of capacities $\mathcal{O}_{\eta}(k^{6})+\mathcal{O}_{\eta}(k^{3r+1})$.
@@
We obtain that $G^{2}$ has a $(\delta,\gamma)$-protrusion decomposition, where $\delta=\mathcal{O}_{\eta}(k^{5})+\mathcal{O}_{\eta}(k^{3r})$ and $\gamma=\mathcal{O}(\eta/\epsilon)$. Let $(G^{3},k_{3})$ be the instance obtained from Lemma 4.6 on input $(G^{2},k_{2})$. Then the number of vertices of $G^{3}$ is $\mathcal{O}_{\mathcal{F}}(\delta)$. Also, if $G^{3}$ has a $k$-sized solution then the treewidth of $G^{3}$ is at most $k+\eta$ and therefore the number of edges of $G^{3}$ is at most $\mathcal{O}_{\mathcal{F}}(k\cdot\delta)$. Finally run the $\beta$-kernelization algorithm for the problem $\mathcal{F}$-Deletion on the instance $(G^{3},k_{3})$. Let $S_{3}$ be a $\beta$-approximate solution for the problem $\mathcal{F}$-Deletion on the instance $(G^{3},k_{3})$.